Top Company holds 90 percent of Bottom Company’s common stock. In the current year, Top reports sales of $908,000 and cost of goods sold of $681,000. For this same period, Bottom has sales of $379,000 and cost of goods sold of $303,200. During the current year, Bottom sold merchandise to Top for $183,000. The parent still possesses 25 percent of this inventory at the current year-end. Bottom had established the transfer price based on its normal gross profit rate. What are the consolidated sales and cost of goods sold?

A)$1,066,100 and $993,350.

B)$1,104,000 and $816,450.

C)$1,104,000 and $846,950.

D)$1,104,000 and $810,350.
Answer #1

Answer:d (1104000 and 810350)

Explanation

Top company :

Sales= 908000

Cost of good sold =681000

Bottom company:

sales= 379000

Cost of good sold 303200

Top sold merchandise to bottom for= 183000

We will compute the cost of good sold

Gross profit= 379000-303200=75800

=75800/379000= 20%

Intra entity gross profit= 183000*25%= 45750*20%= 9150

Top COGS 681000
Bottom COGS 303200
Total COGS 984200
Intra entity transfer (183000)
Intra entity GP 9150
Consolidated COGS 810350

We will compute consolidated sales

Consolidated sales =(top sales + bottom sales -intra entity transfer)

= (908000+379000-183000)= 1104000